the demand curve for a swivel chair is given by p = 4000(3-q) dollars per swivel chair, where p is the price and q is the quantity, in thousands of swivel chair, demanded at that price. what quantity will be demanded if the price per swivel chair is $256.60?


1
Expert's answer
2021-11-03T17:10:59-0400

"\text{The demand curve is given by $p = 4000(3-q)$, when p = 256}\\
\implies 256 = 4000(3-q)\\
\implies q = 2.936"